Northwood Company manufactures basketballs. The company has a ball that sells for $25. At present, the ball is manufactured ineed answers to part 1,5,6A, and 6B
Answer #1

Solution 1:

Contribution margin ratio = Contribution margin / sales = $300,000 / $750,000 = 40%

Contribution margin per unit = $25 - $15 = $10 per unit

Breakeven sales units = Fixed cost / contribution margin per unit = $210,000 / 10 = 21000 units

Degree of operating leverage = Contribution margin / Net operating income = $300,000 / $90,000 = 3.33

Solution 2:

New variable cost per unit = $15 + $3 = $18 per ball

new contribution margin per unit = $25 - $18 = $7 per unit

New contribution margin ratio = $7 / $25 =28%

New breakeven point in balls = $210,000 / $7 = 30000 units

Solution 3:

Nos of balls to be sold to earn target income = (Fixed cost + Target profit) / contribution margin per unit

= ($210,000 + $90,000) / $7 = 42857 units

Solution 4:

Variable cost per unit = $18 per unit

Required contribution margin ratio = 40%

required variable cost ratio = 60%

New selling price per unit = $18 / 60% = $30 per unit

Solution 5:

New variable cost per unit = $15 * 60% = $9 per unit

New contribution margin per unit = 25- $9 = $16 per unit

New fixed costs = $210,000*2 = $420,000

New CM ratio = $16/$25 = 64%

New breakeven point = $420,000/ $16 = 26250 units

Solution 6a:

Nos of balls to be sold to earn target income = (Fixed cost + Target profit) / contribution margin per unit

= ($420,000 + $90,000) / $16 = 31875 units

Solution 6b:

Northwood Company
Contribution margin income statement
Particulars Amount
Sales (30000*$25) $750,000.00
Variable cost (30000*$9) $270,000.00
Contribution margin $480,000.00
Fixed expenses $420,000.00
Net Operating income $60,000.00
Degree of operating leverage (Contribution / Net Operating income) 8.00